OpenStudy (anonymous):

What is the slope of a line parallel to 4x + y = -1?

OpenStudy (butterflydreamer):

Firstly, re-arrange "4x+y = -1" into the y=mx + b form. This way we can find the slope, where the slope is "m" :)

OpenStudy (anonymous):

k

OpenStudy (butterflydreamer):

Basically re-arrange the equation "4x + y = -1" into y= .... (Just make y the subject) LOL .

OpenStudy (anonymous):

y=-4x-1

OpenStudy (anonymous):

so my answer is -4

OpenStudy (butterflydreamer):

correct! :D

OpenStudy (anonymous):

thanks codsack

OpenStudy (butterflydreamer):

you're welcome lol

OpenStudy (anonymous):

wanna help me with some more?

OpenStudy (butterflydreamer):

i don't mind xD Though i'm going to bed in like 6 minutes

OpenStudy (anonymous):

Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5).

OpenStudy (anonymous):

kyles from florida

OpenStudy (butterflydreamer):

first re-arrange your equation again. "x - 3y = 9 " and make y the subject (change it into y = mx+b form to find the slope). Then once you find the slope, since you are trying to find the equation of a line PERPENDICULAR to "x-3y=9" , you have to remember that : m1 times m2 = -1. (where m1 is the slope of your original equation and m2 is the slope of your perp. line). After this and you have found your slope of the perpendicular line, you need to substitute your point (3,5) and the slope into the point slope formula: y - y1 = m (x - x1) [Where y1 = 5 and x1 = 3 (this is from your point (3,5) ] .

OpenStudy (anonymous):

too much to read im tired lol

OpenStudy (anonymous):

kyle thats ungrateful

OpenStudy (anonymous):

flutter you ya wingspan eyed flutter

OpenStudy (butterflydreamer):

lmao well. i tried helping. eh i'm off to bed. good luck

OpenStudy (anonymous):

is it dark there??